i am thinking of a whole number. if i divide it by 13, then multiply the answer by 12, then subtract 20 and then add 89, i end up with 93. what was my original number
Answer by stanbon(26278)

i am thinking of a whole number. if i divide it by 13, then multiply the answer by 12, then subtract 20 and then add 89, i end up with 93. what was my original number
---------------------
Let the number be "x"
Equation:
(x/13)12 - 20 + 90 = 93
(12/13)x + 70 = 93
(12/13)x = 23
x = (13/12)*23
x = 24.92
----------------------
Cheers,
Stan H.